Let Z = f(x,y) = 3x^3 - 5y^2 - 225x + 70y + 23. (i) Find the stationary points of z. (ii) Determine if at these points the function is at arelative maximum, relative minimum, infixion point, or saddle point.
1
Expert's answer
2020-11-09T10:33:06-0500

Z=f(x,y)=3x2-5y2-225x+70y+23

"\tfrac{\partial~f(x,y)}{\partial~x} =9x^{2} -225 =0"

xo ="\sqrt{\dfrac{225}{9}}" =5

"\tfrac{\partial~f(x,y)}{\partial~y} =-10 y +70 =0"

y=-7

D="\begin{vmatrix}
 \tfrac{\partial{^2}~f(x_o,y_o)}{\partial~x^{2}} & \tfrac{\partial{^2}~f(x_o,y_o)}{\partial~y{\partial~x}} \\
 \tfrac{\partial{^2}~f(x_o,y_o)}{\partial~x~\partial~y} & \tfrac{\partial{^2}~f(x_o,y_o)}{\partial~y^{2}}
\end{vmatrix}"

D="\begin{vmatrix}
 90 & 0 \\
 0 & -10
\end{vmatrix}"

D=90*-10=-900

D<0,hence the stationary point (x0,y0) =(5,-7) is a saddle point.